A jar contains two red marbles, three blue marbles, and four green marbles. Phi draws one marble from the jar. Kenny draws a marble from those remaining. What is the probability that Phi draws a green marble and Kenny draws a blue marble? Express your answer as a common fraction.
Answer by stanbon(75887)

A jar contains two red marbles, three blue marbles, and four green marbles. Phi draws one marble from the jar. Kenny draws a marble from those remaining. What is the probability that Phi draws a green marble and Kenny draws a blue marble? Express your answer as a common fraction.
-----------
Phi green: 4/9
Ken blue: 3/8
----
P(phi green AND ken blue) = (4/9)(3/8) = 1/6
====================================================
Cheers,
Stan H.